# Text-to-Audio Generation using Instruction-Guided Latent Diffusion Model

### v3

@@ 5.2. Baselines
Model #Parameters Objective Subjective
FAD $\downarrow$ KL $\downarrow$ IS $\uparrow$ CLAP $\uparrow$ OVL $\uparrow$ REL $\uparrow$
AudioLDM-M-Full-FT $416$M $2.57$ $1.26$ $8.34$ $0.43$- --
AudioLDM-L-Full $739$M $4.18$ $1.76$ $7.76$ $0.43$- --
AudioLDM 2-Full $346$M $2.18$ $1.62$ $6.92$ $0.43$- --
AudioLDM 2-Full-Large $712$M 2.11 $1.54$ $8.29$ $0.44$ $3.56$ $3.19$
Tango-full-FT $866$M $2.51$ $1.15$ $7.87$ $0.54$ $3.81$ $3.77$
Tango 2 $866$M $2.69$ 1.12 9.09 0.57 3.99 4.07
   w/o Strategy 2 & 3 $866$M $2.64$ $1.13$ $8.06$ $0.54$- --
   w/o Strategy 1 $866M$ $2.47$ $1.13$ $8.58$ $0.56$- --
   w/o Strategy 2 $866$M $2.28$ 1.12 $8.38$ $0.55$- --
   w/o Strategy 3 $866$M $2.46$ $1.13$ $8.63$ $0.56$- --
@@ 5.3. Evaluation Metrics
Objective Metrics
We evaluated the text-to-audio generation using the standard Frechet Audio Distance (FAD), KL divergence, Inception score (IS), and CLAP score (Liu et al., 2023a). FAD is adapted from Frechet Inception Distance (FID) and measures the distribution-level gap between generated and reference audio samples. KL divergence is an instance-level reference-dependent metric that measures the divergence between the acoustic event posteriors of the ground truth and the generated audio sample. FAD and KL are computed using PANN, an audio-event tagger. IS evaluates the specificity and coverage of a set of samples, not needing reference audios. IS is inversely proportional to the entropy of the instance posteriors and directly proportional to the entropy of the marginal posteriors. CLAP score is defined as the cosine similarity between the CLAP encodings of an audio and its textual description. We borrowed the AudioLDM evaluation toolkit (Liu et al., 2023a) for the computation of FAD, IS, and KL scores.
@@
Objective Subjective Objective Subjective
@@
Objective Subjective Objective Subjective

### v4

@@ 5.2. Baselines
Model #Parameters Objective – Holistic Objective – Temporal Subjective
FAD $\downarrow$ KL $\downarrow$ IS $\uparrow$ CLAP $\uparrow$ OER $\downarrow$ DUR $\downarrow$ FREQ $\downarrow$ TIME $\uparrow$ OVL $\uparrow$ REL $\uparrow$
AudioLDM-M-Full-FT $416$M $2.57$ $1.26$ $8.34$ $0.43$- -- -- -- -- --
AudioLDM-L-Full $739$M $4.18$ $1.76$ $7.76$ $0.43$- -- -- -- -- --
AudioLDM 2-Full $346$M $2.18$ $1.62$ $6.92$ $0.43$- -- -- -- -- --
AudioLDM 2-Full-Large $712$M 2.11 $1.54$ $8.29$ $0.44$- -- -- -- $3.56$ $3.19$
Tango-full-FT $866$M $2.51$ $1.15$ $7.87$ $0.54$ 0.882 3.535 1.611 0.577 $3.81$ $3.77$
Tango 2 $866$M $2.69$ 1.12 9.09 0.57 0.87 3.586 1.548 0.61 3.99 4.07
   w/o Strategy 2 & 3 $866$M $2.64$ $1.13$ $8.06$ $0.54$- -- -- -- -- --
   w/o Strategy 1 $866M$ $2.47$ $1.13$ $8.58$ $0.56$- -- -- -- -- --
   w/o Strategy 2 $866$M $2.28$ 1.12 $8.38$ $0.55$- -- -- -- -- --
   w/o Strategy 3 $866$M $2.46$ $1.13$ $8.63$ $0.56$ $0.88$ $3.63$ $1.577$ $0.588$- --
@@ 5.3. Evaluation Metrics
Holistic Objective Metrics
We evaluate the generated audio samples in a holistic fashion using the standard Frechet Audio Distance (FAD), KL divergence, Inception score (IS), and CLAP score (Liu et al., 2023a). FAD is adapted from Frechet Inception Distance (FID) and measures the distribution-level gap between generated and reference audio samples. KL divergence is an instance-level reference-dependent metric that measures the divergence between the acoustic event posteriors of the ground truth and the generated audio sample. FAD and KL are computed using PANN, an audio-event tagger. IS evaluates the specificity and coverage of a set of samples, not needing reference audios. IS is inversely proportional to the entropy of the instance posteriors and directly proportional to the entropy of the marginal posteriors. CLAP score is defined as the cosine similarity between the CLAP encodings of an audio and its textual description. We borrowed the AudioLDM evaluation toolkit (Liu et al., 2023a) for the computation of FAD, IS, and KL scores.
Temporal Objective Metrics
To specifically evaluate the temporal controllability of the text-to-audio models, we employ the recently-proposed STEAM (Xie et al., 2024) metrics measured on the AudioTime (Xie et al., 2024) benchmark dataset containing temporally-aligned audio-text pairs. STEAM constitutes four temporal metrics: (i) Ordering Error Rate (OER) – if a pair of events in the generated audio matches their order in the text, (ii) Duration (DUR) / (iii) Frequency (FREQ) – if the duration/frequency of an event in the generated audio matches matches the given text, (iv) Timestamp (TIME) – if the on- and off-set timings of an event in the generated audio match the given text.
@@
Objective – Holistic Subjective Objective – Holistic Subjective
@@
Objective – Holistic Subjective Objective – Holistic Subjective
@@ 5.4. Main Results
A major enhancement in Tango 2 is evident in the temporal objective metrics, as measured by STEAM. With the exception of Duration, Tango 2 shows consistent superiority over Tango across all other temporal measurements. The implementation of Strategy 3 in Audio-alpaca plays a crucial role in temporal data augmentation. Our findings reveal that the absence of this augmentation leads to a decline in Tango 2’s temporal performance, thus highlighting the effectiveness of Strategy 3-based data augmentation technique.
@@ 6. Conclusion
Acknowledgements
This research is supported by the Ministry of Education, Singapore, under its AcRF Tier-2 grant (Project no. T2MOE2008, and Grantor reference no. MOE-T2EP20220-0017).